methane you need two moles of oxygen gas for every mole of methane. If you have five moles of methane When you produce two moles of carbon dioxide and four moles of water. Even though methane is available, you do not have enough oxy Thus you can conclude that oxygen is the limiting reactant Part D Find the number of moles of water that can be formed if you have 154 mol of hydrogen gas and 72 mol of oxygen gas Express your answer with the appropriate units. Hints Value moles Submit Give Up Incorrect; Try Again; 2 attempts remaining 1 pt 8. Net lonic Equations to search
Answer #1

Answer:

2 H2 + O2\rightarrow 2 H2O

figure out the limiting reagent

154 mol H2\times ( 2 mol H2O / 2 mol H2 )= 154 mol H2O

72 mol O2\times ( 2 mol H2O / mol O2 ) = 144 mol H2O

so the maximum you can make it 144 mol H2O because O2 is the limiting reagennt.